a. Find the present values of the following cash flow streams at 8% compounded annually.

0-------------1----------------2--------------3-------------4-------------5

stream a $0------------$100--------$400--------$400---------$400---------$300

stream b $0-----------$300---------$400---------$400--------$400--------$100


1
Expert's answer
2021-05-04T12:54:18-0400

a =


Present value = (100/1.08^1)+(400/1.08^2)+(400/1.08^3)+(400/1.08^4)+(300/1.08^5) =


92.59+342.92+366.06+289.70+204.17


= $1295.4




Stream b


Present value (300/1.08^1)+(400/1.08^2)+(400/1.08^3)+(400/1.08^4)+(100/1.08^5) =


277.78+342.92+366.06+289.70+68.06


=$1344.52